worldlychicken:

Pizza Time has collected data about customer pizza topping preferences. They have calculated that P(pepperoni) = 0.7, P(olives) = 0.6, and P(pepperoni or olives) = 0.8. Determine the P(pepperoni and olives).

ratnakermehta20:

We are given that, P(p)=0.7, P(o)=0.6, and, P(p⋃o)=0.8. Required Probability=P(p⋂o). Recall that, P(p⋃o)=P(p)+P(o)-P(p⋂o). Substituting the given values, we get, 0.8=0.7+0.6-P(p⋂o). ∴ P(p⋂o)=0.7+0.6-0.8=0.5.